OpenStudy (zeig_101):

For the function f(x)=(7-8x)^2, find f^-1. Determine whether f^-1 is a function.

OpenStudy (zeig_101):

Little help guys, I'm not sure how to do things like this.

OpenStudy (anonymous):

what exactly is the questoin asking

OpenStudy (zeig_101):

It's asking to solve that equation for f^-1, and determine whether it's a function or not.

OpenStudy (anonymous):

the inverse is not a function because \[f(x)=(7-8x)^2\] is not a one to one function

OpenStudy (zeig_101):

okay. Can you explain how to find that out?

OpenStudy (anonymous):

the clue is the square

OpenStudy (anonymous):

if you graph the function you will get a parabola, which fails the "vertical line test" for sure

OpenStudy (anonymous):

oops i meant "horizontal line test"

OpenStudy (zeig_101):

Okay. ALSO, do you know how to solve it to find f^-1?

OpenStudy (anonymous):

alternatively you can try to solve \[x=(7-8y)^2\] for \(y\) if you do it carefully you will see that the inverse in not a function

OpenStudy (anonymous):

most common mistake is to solve like this \[x=(7-8y)^2\\ \sqrt{x}=7-8y\\ sqrt{x}-7=-8y\\ \frac{\sqrt{x}-7}{-8}=y\] but that is wrong the first step should be \[\pm\sqrt{x}=7-8y\\ \pm\sqrt{x}-7=-8y\\ \frac{\pm\sqrt{x}-7}{-8}=y\]

OpenStudy (anonymous):

and it is the \(\pm\) that shows the inverse is not a function

OpenStudy (zeig_101):

So f^-1(x)=(7+-(sqrt x))/8?

OpenStudy (anonymous):

yes

OpenStudy (zeig_101):

okay, and it's not a function?

OpenStudy (anonymous):

although it is an abuse of notation to write \[f^{-1}(x)=\frac{\pm\sqrt{x}-7}{8}\] because when you write \(f^{-1}\) you are saying it is a function, even though it is not yes, it is not a function

OpenStudy (zeig_101):

Can you help me with another question?

OpenStudy (anonymous):

go ahead and post, i will look at it

OpenStudy (zeig_101):

g(x)=14/(x+3); Find (g^-1*o*g)(4)